# On local Galois deformation rings: generalised tori

### v1

@@ 8. Rank calculations
If $N$ is finitely generated over $\mathbb{Z}$ then the statement is proved in [4, Corollary 2, p. 105]. The same argument carries over if $N$ is finitely generated over $\mathbb{Z}_{p}$: the cohomology groups are finitely generated $\mathbb{Z}_{p}$-modules, since the complex computing cohomology consists of finitely generated $\mathbb{Z}_{p}$-modules, moreover they are annihilated by the order of $\Delta$. ∎
@@
We choose $V$ to be the image of a $p$-adic exponential function defined on $\mathfrak{p}_{E}^{n}$ for some large enough $n\geq 1$. We then have an isomorphism $\mathfrak{p}_{E}^{n}\otimes M\cong V\otimes M$ of $\Delta$-modules. Since $\mathfrak{p}_{E}^{n}\cong\mathcal{O}_{E}$ is isomorphic to $\mathcal{O}_{F}[\Delta]$ as $\mathbb{Z}_{p}[\Delta]$-module, see the proof of [13, Section 1.4], $\mathcal{O}_{E}\otimes M$ is free and thus $(\mathcal{O}_{E}\otimes M)_{\Delta}\cong\mathcal{O}_{F}\otimes M$. Thus $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{O}_{E}^{\times})^{\wedge,p}\otimes M)_{\Delta}=[F:\mathbb{Q}_{p}]\cdot\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M$ and the assertion follows. ∎
@@
We have $\mathcal{A}^{\wedge,p}\cong\mathcal{A}\otimes\mathbb{Z}_{p}$ and $(\mathcal{A}^{\Delta})^{\wedge,p}\cong\mathcal{A}^{\Delta}\otimes\mathbb{Z}_{p}$ by [14, Tag 00MA]. We may express $\mathcal{A}^{\Delta}$ as the kernel of
@@
Since $M$ is a free $\mathbb{Z}$-module of finite rank and $(E^{\times})^{\wedge,p}\cong\Gamma^{\mathrm{ab},p}_{E}$ we $(E^{\times}\otimes M)^{\wedge,p}\cong(E^{\times})^{\wedge,p}\otimes M\cong\Gamma^{\mathrm{ab},p}_{E}\otimes M$. Since $I_{\Delta}$ is a free $\mathbb{Z}$-module $(I_{\Delta}\otimes M)^{\wedge,p}\cong I_{\Delta}\otimes M^{\wedge,p}$, so that we obtain an exact sequence of $\mathbb{Z}_{p}[\Delta]$-modules:
@@ 9. Galois deformations
We will apply the machinery developed in the previous sections in an arithmetic situation. Let $F$ be a finite extension of $\mathbb{Q}_{p}$. We fix an algebraic closure $\overline{F}$ and let $\Gamma_{F}:=\operatorname{Gal}(\overline{F}/F)$. Let $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ be a continuous representation, where $G$ is a generalised torus over $\mathcal{O}$ and let $\overline{\Theta}$ be the $G$-pseudocharacter associated to $\overline{\rho}$. Let $\Gamma_{E}$ be the kernel of the composition $\Gamma_{F}\overset{\overline{\rho}}{\longrightarrow}G(k)\rightarrow(G/G^{0})(k)$ and let $\Delta:=\operatorname{Gal}(E/F)$. Let $\pi:\Gamma_{F}\rightarrow\Delta$ be the projection map.
@@
corresponding to the fundamental class $[u_{E/F}]\in H^{2}(\Delta,E^{\times})$ by [15, (1.2)]. Let $\widehat{W}_{E/F}$ be the profinite completion of $W_{E/F}$. The Artin map $\operatorname{Art}_{E}:E^{\times}\rightarrow\Gamma_{E}^{\mathrm{ab}}$ induces an isomorphism between the profinite completion of $E^{\times}$ and $\Gamma_{E}^{\mathrm{ab}}$. This yields a natural isomorphism
@@
Every continuous representation $\rho:\Gamma_{F}\rightarrow G(A)$, where $A$ is a finite discrete $\mathcal{O}$-algebra, factors through the quotient $\Gamma_{F}\twoheadrightarrow\widehat{W}_{E/F}$.
@@
If $L^{\prime}$ is a finite extension of $L$ with the ring of integers $\mathcal{O}^{\prime}$ and residue field $k^{\prime}$ then the functor $D^{\mathrm{ps}}_{\overline{\Theta}_{k^{\prime}}}:\mathfrak{A}_{\mathcal{O}^{\prime}}\rightarrow\operatorname{Set}$ is pro-representable by $R^{\mathrm{ps}}_{\overline{\Theta}}\otimes_{\mathcal{O}}\mathcal{O}^{\prime}$ and analogous statements hold for $X^{\mathrm{gen}}_{G,\overline{\Theta}_{k^{\prime}}}$ and $D^{\square}_{\overline{\rho}_{k^{\prime}}}$. Thus it is enough to prove the statement for $\mathcal{O}^{\prime}=\mathcal{O}$, while allowing to replace $L$ by a finite extension.
@@
Lemma 7.10 and (65) imply that $R^{\mathrm{ps}}_{\overline{\Theta}}$ is isomorphic to $\mathcal{O}\llbracket N\rrbracket$, where $N$ is the pro-$p$ completion of $(E^{\times}\otimes M)^{\Delta}$. It follows from Lemma 8.7 that $N\cong(\Gamma^{\mathrm{ab}}_{E}\otimes M)^{\Delta}$, which is a finitely generated $\mathbb{Z}_{p}$-module of rank $r$ and torsion subgroup isomorphic to $(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}$ by 8.8. This yields the isomorphisms in part (3). Part (4) follows from Proposition 7.13. Part (5) follows from part (4) and Lemma 7.4 as $E^{\times}/(E^{\times})^{p}$ is finite and the map $\rho\mapsto\Theta_{\rho}$ induces a map of local $\mathcal{O}$-algebras $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$. ∎
@@
Since the map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow\mathcal{O}^{\prime}\otimes_{\mathcal{O}}R^{\mathrm{ps}}_{\overline{\Theta}}$ is faithfully flat, the assertion follows from [14, Tag 06CM] and part (5) of Theorem 9.3. ∎
@@
The map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow A^{\mathrm{gen}}_{G,\overline{\Theta}}$ is formally smooth by the same argument as in Corollary 9.4 using part (4) of Theorem 9.3. Since it is of finite type it is smooth by [14, Tag 00TN]. The map $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}:=\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ is flat and of finite presentation and hence open by [14, Tag 00I1]. By [14, Tag 004Z] it is enough to show the fibres of $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}$ are irreducible.
@@
Lemma 9.6 implies that the scheme $X^{\mathrm{gen}}_{G,\overline{\Theta}}$ coincides with the scheme $X^{\mathrm{gen}}_{G,\bar{\rho}^{\mathrm{ss}}}$ defined in [11] for a generalised reductive group $G$, see [11, Proposition 7.3].
@@
Let $f:\mathcal{P}_{1}\rightarrow\mathcal{P}_{2}$ be a homomorphism of finitely generated $\mathbb{Z}_{p}$-modules and let $R_{i}=k\llbracket\mathcal{P}_{i}\rrbracket$. If $\ker f$ is torsion then $\mathcal{P}_{1}^{\mathrm{tf}}$ is a submodule of $\mathcal{P}_{2}^{\mathrm{tf}}$ and the fibre $k\otimes_{R_{1}^{\mathrm{red}}}R_{2}^{\mathrm{red}}$ is isomorphic to $k\llbracket\mathcal{P}^{\mathrm{tf}}_{2}/\mathcal{P}^{\mathrm{tf}}_{1}\rrbracket$. Since the dimension of the fibre is equal to $d_{2}-d_{1}$, where $d_{i}=\dim_{\mathbb{Q}_{p}}(\mathcal{P}_{i}\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p})$, we deduce from [14, Tag 00R4] that $R_{1}^{\mathrm{red}}\rightarrow R_{2}^{\mathrm{red}}$ is flat.

### v2

@@ 8. Rank calculations
If $N$ is finitely generated over $\mathbb{Z}$ then the statement is proved in [5, Corollary 2, p. 105]. The same argument carries over if $N$ is finitely generated over $\mathbb{Z}_{p}$: the cohomology groups are finitely generated $\mathbb{Z}_{p}$-modules, since the complex computing cohomology consists of finitely generated $\mathbb{Z}_{p}$-modules, moreover they are annihilated by the order of $\Delta$. ∎
@@
We choose $V$ to be the image of a $p$-adic exponential function defined on $\mathfrak{p}_{E}^{n}$ for some large enough $n\geq 1$. We then have an isomorphism $\mathfrak{p}_{E}^{n}\otimes M\cong V\otimes M$ of $\Delta$-modules. Since $\mathfrak{p}_{E}^{n}\cong\mathcal{O}_{E}$ is isomorphic to $\mathcal{O}_{F}[\Delta]$ as $\mathbb{Z}_{p}[\Delta]$-module, see the proof of [14, Section 1.4], $\mathcal{O}_{E}\otimes M$ is free and thus $(\mathcal{O}_{E}\otimes M)_{\Delta}\cong\mathcal{O}_{F}\otimes M$. Thus $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{O}_{E}^{\times})^{\wedge,p}\otimes M)_{\Delta}=[F:\mathbb{Q}_{p}]\cdot\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M$ and the assertion follows. ∎
@@
We have $\mathcal{A}^{\wedge,p}\cong\mathcal{A}\otimes\mathbb{Z}_{p}$ and $(\mathcal{A}^{\Delta})^{\wedge,p}\cong\mathcal{A}^{\Delta}\otimes\mathbb{Z}_{p}$ by [15, Tag 00MA]. We may express $\mathcal{A}^{\Delta}$ as the kernel of
@@
Since $M$ is a free $\mathbb{Z}$-module of finite rank and $(E^{\times})^{\wedge,p}\cong\Gamma^{\mathrm{ab},p}_{E}$ we have
$(E^{\times}\otimes M)^{\wedge,p}\cong(E^{\times})^{\wedge,p}\otimes M\cong\Gamma^{\mathrm{ab},p}_{E}\otimes M.$
Since $I_{\Delta}$ is a free $\mathbb{Z}$-module $(I_{\Delta}\otimes M)^{\wedge,p}\cong I_{\Delta}\otimes M^{\wedge,p}$, so that we obtain an exact sequence of $\mathbb{Z}_{p}[\Delta]$-modules:
@@ 9. Galois deformations
We will apply the machinery developed in the previous sections in an arithmetic situation. Let $F$ be a finite extension of $\mathbb{Q}_{p}$. We fix an algebraic closure $\overline{F}$ and let $\Gamma_{F}:=\operatorname{Gal}(\overline{F}/F)$. Let $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ be a continuous representation, where $G$ is a generalised torus over $\mathcal{O}$ and let $\overline{\Theta}$ be the $G$-pseudocharacter associated to $\overline{\rho}$. Let $\Gamma_{E}$ be the kernel of the composition $\Gamma_{F}\overset{\overline{\rho}}{\longrightarrow}G(k)\rightarrow(G/G^{0})(k)$ and let $\Delta:=\operatorname{Gal}(E/F)$. Let $\pi:\Gamma_{F}\rightarrow\Delta$ and $\Pi:G\rightarrow G/G^{0}$ be the projection maps.
@@
corresponding to the fundamental class $[u_{E/F}]\in H^{2}(\Delta,E^{\times})$ by [16, (1.2)]. Let $\widehat{W}_{E/F}$ be the profinite completion of $W_{E/F}$. The Artin map $\operatorname{Art}_{E}:E^{\times}\rightarrow\Gamma_{E}^{\mathrm{ab}}$ induces an isomorphism between the profinite completion of $E^{\times}$ and $\Gamma_{E}^{\mathrm{ab}}$. This yields a natural isomorphism
@@
Every continuous representation $\rho:\Gamma_{F}\rightarrow G(A)$ satisfying $\Pi\circ\rho=\pi$, where $A$ is a finite discrete $\mathcal{O}$-algebra, factors through the quotient $\Gamma_{F}\twoheadrightarrow\widehat{W}_{E/F}$.
@@
If $L^{\prime}$ is a finite extension of $L$ with the ring of integers $\mathcal{O}^{\prime}$ and residue field $k^{\prime}$ then the functor $D^{\mathrm{ps}}_{\overline{\Theta}_{k^{\prime}}}:\mathfrak{A}_{\mathcal{O}^{\prime}}\rightarrow\operatorname{Set}$ is pro-representable by $R^{\mathrm{ps}}_{\overline{\Theta}}\otimes_{\mathcal{O}}\mathcal{O}^{\prime}$ and analogous statements hold for $X^{\mathrm{gen}}_{G,\overline{\Theta}_{k^{\prime}}}$ and $D^{\square}_{\overline{\rho}_{k^{\prime}}}$. Thus it is enough to prove the statement for $\mathcal{O}^{\prime}=\mathcal{O}$, after replacing $L$ by a finite extension.
@@
Lemma 7.10 and (65) imply that $R^{\mathrm{ps}}_{\overline{\Theta}}$ is isomorphic to $\mathcal{O}\llbracket N\rrbracket$, where $N$ is the pro-$p$ completion of $(E^{\times}\otimes M)^{\Delta}$. It follows from Lemma 8.7 that $N\cong(\Gamma^{\mathrm{ab},p}_{E}\otimes M)^{\Delta}$, which is a finitely generated $\mathbb{Z}_{p}$-module of rank $r$ and torsion subgroup isomorphic to $(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}$ by 8.8. This yields the isomorphisms in part (3). Part (4) follows from Proposition 7.13.
The map $\rho\mapsto\Theta_{\rho}$ induces a map of local $\mathcal{O}$-algebras $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$ and hence a map of $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebras $A^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$. Since $E^{\times}/(E^{\times})^{p}$ is finite, Lemma 7.4 implies that $R^{\square}_{\overline{\rho}}$ is the completion of $\mathcal{O}(\operatorname{Rep}^{W_{E/F}}_{G,\pi})$ with respect to the maximal ideal corresponding to $\overline{\rho}$. By considering the composition $\mathcal{O}(\operatorname{Rep}^{W_{E/F}}_{G,\pi})\rightarrow A^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$ we deduce that that it induces a natural isomorphism between $R^{\square}_{\overline{\rho}}$ and the completion of $A^{\mathrm{gen}}_{G,\overline{\Theta}}$ with respect to the maximal ideal corresponding to $\overline{\rho}$. Part (5) then follows from parts (3) and (4). ∎
@@
Since the map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow\mathcal{O}^{\prime}\otimes_{\mathcal{O}}R^{\mathrm{ps}}_{\overline{\Theta}}$ is faithfully flat, the assertion follows from [15, Tag 06CM] and part (5) of Theorem 9.3. ∎
@@
The map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow A^{\mathrm{gen}}_{G,\overline{\Theta}}$ is formally smooth by the same argument as in Corollary 9.4 using part (4) of Theorem 9.3. Since it is of finite type it is smooth by [15, Tag 00TN]. The map $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}:=\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ is flat and of finite presentation and hence open by [15, Tag 00I1]. By [15, Tag 004Z] it is enough to show the fibres of $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}$ are irreducible.
@@
Lemma 9.6 implies that the scheme $X^{\mathrm{gen}}_{G,\overline{\Theta}}$ coincides with the scheme $X^{\mathrm{gen}}_{G,\bar{\rho}^{\mathrm{ss}}}$ defined in [12] for a generalised reductive group $G$, see [12, Proposition 7.3].
@@
Let $f:\mathcal{P}_{1}\rightarrow\mathcal{P}_{2}$ be a homomorphism of finitely generated $\mathbb{Z}_{p}$-modules and let $R_{i}=k\llbracket\mathcal{P}_{i}\rrbracket$. If $\ker f$ is torsion then $\mathcal{P}_{1}^{\mathrm{tf}}$ is a submodule of $\mathcal{P}_{2}^{\mathrm{tf}}$ and the fibre $k\otimes_{R_{1}^{\mathrm{red}}}R_{2}^{\mathrm{red}}$ is isomorphic to $k\llbracket\mathcal{P}^{\mathrm{tf}}_{2}/\mathcal{P}^{\mathrm{tf}}_{1}\rrbracket$. Since the dimension of the fibre is equal to $d_{2}-d_{1}$, where $d_{i}=\dim_{\mathbb{Q}_{p}}(\mathcal{P}_{i}\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p})$, we deduce from [15, Tag 00R4] that $R_{1}^{\mathrm{red}}\rightarrow R_{2}^{\mathrm{red}}$ is flat.